# Scaling Law with Learning Rate Annealing

### v1

@@ 3.2 LR Annealing Momentum
(a) The different delay steps in the annealing process with different annealing steps (0.1K, 0.5K, 1K and 2K) and the same min learning rate.
(b) The different delay steps in the re-warmup process with different re-warmup steps (0.1K, 0.5K, 1K and 2K) and the same max learning rate.
@@
Figure 5: The delay phenomenon between the learning rate and validation loss. This phenomenon suggests that learning rate annealing (re-warmup) has momentum.
@@
Another key observation is that LR annealing has momentum. For more information about the property of $f(\eta)$, We design one special LRS where the learning rate linearly decreases from $\eta_{max}$ to $\eta_{min}$ then increases. We maintain the slope of the increasing stage (always 5K steps) and change the slope in the decreasing stage. The decreasing steps include 0.1K, 0.5K, 1K, and 2K. Symmetrically, we design another special LRS where the learning rate linearly increases from $\eta_{min}$ to $\eta_{max}$ then decreases. We observe how the validation loss changes with LR increasing or decreasing. The learning rate and loss curves are shown in Fig. 5.
@@
We discover the delay phenomenon between learning rate and validation loss. Firstly, the turning point of validation loss curve is always later than the turning point of learning rate curve in both schedulers. That means that the validation loss still maintains its previous trajectory for a few steps even after the learning rate changes direction. Secondly, the steeper the slope of the learning rate annealing (or re-warmup), the more pronounced the delay between the two turning points. Thirdly, given the same LR slope, the left figure (LR decreasing then increasing) always has a longer delay steps than the right figure (LR decreasing then increasing).
@@
Interestingly, the entire experimental phenomenon is strikingly similar to the physical experiment of a small ball falling along a slope. The larger the angle of the slope, the faster the ball falls. When the ball lands, the accumulated momentum causes the ball to continue sliding for some distance. Based on the delay phenomenon, we pose a hypothesis that $f(\eta)$, the loss drop brought by LR annealing has cumulative historical formation so that the past change of learning rate will affect the following loss curve for a few steps. In summary, learning rate annealing has momentum.
Thus we define $f(\eta)=C\cdot S_{2}$ as follows:
$m_{i}$ $=\lambda\cdot m_{i-1}+(\eta_{i-1}-\eta_{i})$ (6)
$S_{2}$ $=\sum\limits_{i=1}^{s}m_{i}$
$=\sum\limits_{i=1}^{s}\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k}$
@@
Where $m_{i}$ is the LR annealing momentum at step $i$, and $\Delta\eta=\eta_{i-1}-\eta_{i}$ denotes the LR annealing amount. $\lambda$ is the decay factor, which signifies the degree to which historical information is retained. We find that $\lambda$ empirically ranges from $0.99$ to $0.999$. In fact, if LR annealing had no momentum, $\lambda=0$ would induce that $f(\eta)=C\cdot\eta_{s}$, which is exactly the naive format mentioned above. Note that $S_{2}$ is not only applicable to LR annealing ($S_{2}>0$), but also to LR re-warmup ($S_{2}<0$). This means that our equation can be also applied in continual pre-training, where LR re-warmup serves as an important factor for better outcomes. More intuitively, the definition of $S_{2}$ can be visualized in Fig. 1, as the weighted sum of the areas of the small blue squares.
3.3 Final Format
Definition.
@@ 3.3 Final Format
Given the same training and validation dataset, the same model size, the same training hyper-parameters such as max learning rate $\eta_{max}$ and batch size, the language modeling loss at step $s$ in a full loss curve empirically follow as $L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$, where $S_{1}=\sum\limits_{i=1}^{s}\eta_{i}$ and $S_{2}$ are defined as Eq. 6. $L_{0}$, $A$, $C$, $\alpha$ are all undetermined positive constants to be fitted.
It is feasible to apply the formulation to universally describe the loss of each step across diverse learning rate schedulers. This formulation supports validation loss prediction in a more complicated LRS of longer total steps, fitted from a simpler LRS of shorter total steps.
@@
Notice that $\frac{\partial L}{\partial S_{1}}<0$ and $\frac{\partial L}{\partial S_{2}}<0$ always hold true, which means both the increase of S1 and S2 can help to reduce the loss. However, as shown intuitively in Fig. 1, there exists a delicate balance between $S_{1}$ and $S_{2}$. Once the learning rate begins to anneal and $S_{2}$ starts to increase, the forward area $S_{1}$ of incoming steps starts to diminish. Our equation aptly describes this delicate balance. In Sec. 4, we elaborate in detail on this issue.
@@ 3.4 Experiments
Different warmup steps can result in different loss curves in training from scratch. During the warmup stage, neural networks are prone to random optimization, resulting in unpredictable outcomes in the very beginning stage (Hestness et al., 2017). One evidence is about the high gradient norm in the LR warmup stage. Fig. 4 shows that at the very beginning of training, the gradient norm initially converges from a relatively high value to a lower one, exhibiting a pattern distinct from the subsequent steps. From another perspective, LR warmup is also necessary. Our pilot experiment (refer to Appendix A) shows that warmup indeed significantly accelerates convergence, a finding also noted by Liu et al. (2020); Kosson et al. (2024). Consequently, in all experiments conducted for this paper, we employ a 500-step learning rate linear warmup to reach the max LR, where $S_{1}$ and $S_{2}$ are computed yet as constant max LR in the warmup stage  [^3] Note LR warmup in training from scratch is different from LR re-warmup in continual training, where we do not regard re-warmup steps as a hyper-parameter and will show how to apply our equation to find optimal re-warmup steps in Sec. 4..
@@
In this paper, we adopt standard experimental setups in large language model pre-training. The training dataset is Fineweb (Penedo et al., 2024) and the validation dataset is RedPajama-CC (Computer, 2023). We train a 594M non-embedding parameters LLAMA architecture-like model (Meta, 2024) from scratch. We use AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The weight decay is set as $0.1$ and gradient clip is set as $1.0$. We set maximal learning rate as $2\times 10^{-4}$ and batch size as $4$M tokens. We use the tokenizer of LLAMA-3 (Meta, 2024). We adopt the decay factor of learning rate annealing $\lambda=0.999$ and we discuss the impact of $\lambda$ in Sec. 6.
To validate that our formulation works across different experimental settings, we also conduct our experiments on another experimental setups. Please refer to Appendix B for full experimental setups.
@@
Given a learning rate scheduler, we can easily compute out $S_{1}$ and $S_{2}$ of each step in advance. To estimate $(L_{0},A,C,\alpha)$, we adopt a similar fitting method as chinchilla scaling law (Hoffmann et al., 2022). Specifically, we minimize the Huber loss (Huber, 1964) between the predicted and the observed log loss using the L-BFGS algorithm (Nocedal, 1980):
$\min_{L_{0},A,C,\alpha}\quad\sum_{\mathrm{Step~{}}i}\mathrm{Huber}_{\delta}\left(\log\hat{L}(i)-\log L(i)\right)$ (7)
We implement this by the utilization of minimize in scipy library. Huber loss is to enhance to robustness of the fitting results and we set $\delta$ of Huber loss as $10^{-3}$. We mitigate the potential issue of local minima of fitting by choosing the optimal fit from a range of initial conditions. Note that in practice, we can also fit the full loss curves using multiple LRS with a single tuple of $(L_{0},A,C,\alpha)$. In this situation, we sum the Huber losses in Eq. 7 of all fitted LRS.
@@
We employ the full loss curves from constant and cosine LRS, both containing 20K total steps, as our fitting data points (refer to Fig. 2). We then proceed to predict the full loss curves for various unseen LRS with 60K total steps (refer to Fig. 3). The coefficient of determination ($R^{2}$) in the fitting process is larger than $0.999$, nearly perfect, which underscores the robust capability of our equation to fit loss curves across diverse LRS using a single parameter tuple.
As for prediction shown in Fig. 3, our equation demonstrates broad applicability and robust generalization across five distinct types of LRS, including four previously unseen LRS. This is evidenced by the low mean prediction error, which is approximately 0.2%. Remarkably, our equation is capable to accurately predict losses even in a very complicated LRS with multiple LR re-warmup stages, as illustrated in Fig. 3(e), despite the absence of any LR re-warmup stage in the fitting LRS. We also fit and predict on the curves adopting another experimental setups (shown in Appendix C), and the results are similarly good, which proves that Eq. 1 holds across different experimental setups.
@@ 3.5 Extension to Model Size Scaling
Loss Drop in Annealing Stage Scales with Model Size $N$.
First, we research whether the model size $N$ impacts on the amount of loss drop in the annealing stages. We compare the difference of final losses between constant LRS and WSD LRS (10% cosine annealing to $\eta_{min}=0$), thus to get the loss gap brought by LR annealing. We conduct this experiment on different total steps and different model sizes. The experimental results are shown in Fig. 6(a). It suggests that the loss drop brought by LR annealing is scaling with both annealing steps and model size, which means that the annealing area $S_{2}$ in our equation gets larger when model size $N$ becomes large. We suppose there is a simple relationship of $S_{2}\propto N^{\gamma}$ where $\gamma$ is a positive constant to be fitted.
(a) The loss drop brought by LR annealing of different model sizes. The dashed lines represent the trend over steps. It shows that the loss gap brought by LR annealing also scales with data size and model size.
(b) Curve fitting on cosine LRS (60K steps to $\eta_{min}=0.1\cdot\eta_{max}$) of many model sizes using our scaling law extended to model size. By substituting specific values of $N$, the obtained equation simplifies to Eq. 1.
Figure 6: The loss drop brought by LR annealing (on the left) and the $N$-extended full loss curve fitting (on the right). It suggests that $S_{2}\propto N^{\gamma}$ is reasonable from both actual experimental phenomena and curve fitting.
The Format with Model Size.
According to the experiments and analysis above, based on the typical scaling law (Eq. 2), we can extend our proposed equation to model size scaling as follows:
$L(s,N)=L_{0}+A\cdot S_{1}^{-\alpha}+B\cdot N^{-\beta}-C\cdot S_{2}\cdot N^{\gamma}$ (8)
where $N$ denotes the number of non-embedding model parameters (M), and $B$, $\beta$, $\gamma$ are $N$-related undetermined positive constants.
Fitting with Model Size.
We utilize the Eq. 8 to fit full loss curves of different model sizes. The fitting result is shown in Fig. 6(b), where $R^{2}>0.999$ is very close to $1$, suggesting that the loss curves of different model sizes indeed follow our proposed $N$-extended equation. We also fit the curves adopting another experimental setups (shown in Appendix D) and get similar results, proving that Eq. 8 holds across different experimental setups.
@@ 4 Takeaways: Experimental Findings Verification and Explanation
Our derived equation describes the training dynamics of language models. In this section, we apply the equation to give theoretical verification and explanation for many existing experimental findings. These key insights also serve as a guide for researchers to select critical LRS in advance by prediction via our equation, with little computational cost. An interesting summary is that
@@
We adopt our equation to help researchers understand why loss drops more sharply when LR anneals, which has been widely observed in many previous studies. We substitute the fitted parameters (see Fig. 2) to our equation as an instance. We draw how the $S_{1}$-item ($A\cdot S_{1}^{-\alpha}$) and the negative $S_{2}$-item ($-C\cdot S_{2}$) impacts the loss along with a WSD scheduler. Fig. 7 suggests that starting from annealing stage, negative $S_{2}$-item has a much more significant impact on the overall loss than $S_{1}$-item, which makes loss drop more sharply compared with the stable LR stage. In conclusion, LR annealing brings out quick increase of the annealing area, resulting in a drastic decrease in validation loss.
Figure 7: How S1-item and negative $S_{2}$-item changes in a WSD scheduler. Gray area means the amount of loss drop brought by $S_{1}$ and $S_{2}$ in annealing stage.
4.2 It verifies and explains when we use cosine LRS, we should set the cosine cycle length $T$ as the total steps $S$, and set min LR as 00 to get the optimal loss.
Figure 8: The predicted loss curves of different cycle length $T$ and min LR in cosine LRS, well aligned with previous studies.
Many papers have found that in LLM pre-training using cosine LRS, setting the cosine cycle length $T$ as the total steps $S$, and setting min LR as nearly 0 (rather than 10% max LR) can lead to the optimal loss (Hoffmann et al., 2022; Hu et al., 2024; Parmar et al., 2024). Actually, the settings above have been a factual standard in LLM pre-training using cosine LRS. We theoretically verify and explain the finding by our equation, presented as Fig. 8. The prediction curve convincingly demonstrates that the loss curve, with the configuration $T=S$ and a min LR of 0, indeed achieves the optimal loss in the end.
Moreover, our equation gives a quite intuitive explanation. $T>S$ leads to an incomplete annealing while $T<S$ leads to a small forward area due to early annealing. Thus, it is optimal to set $T$ as the total steps $S$. Similarly, setting min LR as 00 can make larger annealing amount and thus larger annealing area $S_{2}$, which enables lower final loss.
@@
(a) The predicted loss curve of 10K steps.
(b) The predicted loss curve of 100K steps
@@
Figure 9: Comparison of constant and cosine in different steps.
In the experiments, we find that if we set small total steps, the final loss of constant LRS could be even lower than cosine LRS, and vice versa. Refer to the ground-truth loss in Fig. 2 (20K steps), and the ground-truth loss in Fig. 3(a)3(b) (60K steps). To validate this phenomenon, we use our equation to draw the prediction loss curve of 10K total steps and 100K total steps in Fig. 9. It shows that our proposed equation can verify well that the better LRS changes over the total steps. Moreover, Fig. 9(c) shows the predicted final loss of different total steps using constant and cosine LRS. It further convincingly suggests that constant LRS indeed gets a lower loss if setting small total steps, but the scaling slope is smaller than cosine LRS’s, resulting in higher loss in more steps.
From a more essential and comprehensive perspective, $|\frac{\partial L}{\partial S_{1}}|$ is a power-law decreasing function while $|\frac{\partial L}{\partial S_{2}}|$ is stable over training steps. In the early stages, $|\frac{\partial L}{\partial S_{1}}|$ is large when $S_{1}$ is small, thus increasing $S_{1}$ by maintaining large LR (e.g. constant LRS) in the early stages can greatly help reduce the loss. That is, $S_{1}$ plays a dominant role over $S_{2}$. In the later stages, $|\frac{\partial L}{\partial S_{1}}|$ is much smaller when $S_{1}$ becomes large, thus increasing $S_{1}$ in the later stages does not significantly help reduce the loss. That is, $S_{2}$ plays a dominant role over $S_{1}$. At this stage, It is time to start LR annealing to increase $S_{2}$. Interestingly, this perspective aligns directly with the idea of WSD LRS (Hu et al., 2024): In the early stages, the neural network is exploring globally and it is a suitable time to use a larger LR; In the later stages, the neural network is exploring locally and it is a suitable time to use a smaller LR. We will delve further into WSD LRS in the following subsections.

### v2

@@ 3.2 LR Annealing Momentum
(a) Different delay steps in the annealing process associated with different annealing steps (0.1K, 0.5K, 1K and 2K).
(b) Different delay steps in the re-warmup process associated with different re-warmup steps (0.1K, 0.5K, 1K and 2K).
@@
Figure 5: The delay phenomenon between the LR and validation loss curves. This phenomenon suggests that LR annealing (re-warmup) has momentum.
@@
Another key observation is that LR annealing has momentum. To refine the formulation of $f(\eta)$, we design a special LRS where the LR decreases linearly from $\eta_{max}$ to $\eta_{min}$ and then increases. The increasing stage always has a fixed slope, reaching the maximum value in 5K steps, while the slope of the decreasing stage is varied, with durations of 0.1K, 0.5K, 1K, and 2K. Symmetrically, we design another LRS where the LR increases linearly from $\eta_{min}$ to $\eta_{max}$ and then decreases. Fig. 5 shows the corresponding LR and loss curves.
@@
We observe a delay phenomenon between the LR and the validation loss. Firstly, the turning point of the validation loss curve consistently lags behind the turning point of the LR curve, indicating that the validation loss continuous along its previous trajectory for some steps even after the LR changes direction. Secondly, the steeper the slope of the LR annealing (or re-warmup), the more pronounced the delay phenomenon becomes. Thirdly, given the same LR slope, the left figure (where LR decreases then increases) consistently shows a longer delay compared to the right figure (where LR increases then decreases). We discuss some possible root reasons of delay phenomenon in Sec. 6.2.
@@
Interestingly, this phenomenon closely resembles the physical experiment of a small ball rolling down a slope. The steeper the slope, the faster the ball accelerates. When the ball lands, the accumulated momentum causes the ball to slide further. Inspired by this delay phenomenon, we hypothesize that $f(\eta)$, the loss reduction induced by LR annealing, has cumulative historical formation so that the past change of learning rate will affect the following loss curve for a few steps. In summary, learning rate annealing exhibits momentum. To capture this, we define $f(\eta)=C\cdot S_{2}$, where $S_{2}$ is calculated as:
$m_{i}$ $=\lambda\cdot m_{i-1}+(\eta_{i-1}-\eta_{i}),$ (5)
$S_{2}$ $=\sum\limits_{i=1}^{s}m_{i}=\sum\limits_{i=1}^{s}\sum\limits_{k=1}^{i}(\eta_{k-1}-\eta_{k})\cdot\lambda^{i-k},$
@@
where $m_{i}$ is the LR annealing momentum at step $i$ ($m_{1}=0$), and $\Delta\eta=\eta_{i-1}-\eta_{i}$ denotes the LR annealing amount at step $i$. $\lambda$ is the decay factor that signifies how much historical information is retained. We find that $\lambda$ values between $0.99$ and $0.999$ generally works well. In contrast, $\lambda=0$ implies no momentum effect, reducing $f(\eta)$ to $C\cdot\eta_{s}$, which degenerate to the initial form mentioned above. Note that $S_{2}$ applies not only to LR annealing ($S_{2}>0$), but also to LR re-warmup ($S_{2}<0$). This means that our equation is applicable to scenarios like continual pre-training, where LR re-warmup serves as an important factor for better outcomes (see Sec. 4). More intuitively, the definition of $S_{2}$ can be visualized in Fig. 1, as the weighted sum of blue grid areas.
3.3 Final Formulation
Scaling Law with LR Annealing.
@@ 3.3 Final Formulation
Given the same training and validation dataset, the same model size, the same training hyper-parameters such as warmup steps, max learning rate $\eta_{max}$ and batch size, the language modeling loss at training step $s$ empirically follows the equation $L(s)=L_{0}+A\cdot S_{1}^{-\alpha}-C\cdot S_{2}$, where $S_{1}$ and $S_{2}$ are defined in Eq. 1. $L_{0}$, $A$, $C$, $\alpha$ are positive constants.
Our formulation describes the loss of each training step across different LRS. It allows fitting based on a simpler LRS with shorter training steps and enables the prediction of validation losses for more complex LRS with longer training steps.
Figure 6: Loss surface of language models as a slide after simplification. Optimization direction could be decomposed into two directions: power-law scaling direction ($S_{1}$, sliding down) and annealing direction ($S_{2}$, inner height of the slide).
Only One Extra Parameter.
Notably, we add only one parameter, the coefficient of $S_{2}$ term $C$, compared to Chinchilla scaling law (Hoffmann et al., 2022). That is, we utilize the fewest extra parameters but model the essential training dynamics in LR annealing to the greatest extent.
Loss Surface as a Slide.
To better understand our formulation, we view the loss surface of language models as a slide in Fig. 6. The optimization process can be seen as sliding down the slide according to the power-law scaling ( orange line), while oscillating on the inner wall ( blue dashed line). When the learning rate anneals ( red line), the amplitude of the oscillation decreases, resulting in a reduction in loss.
@@
Note that in Eq. 1, $\frac{\partial L}{\partial S_{1}}<0$ and $\frac{\partial L}{\partial S_{2}}<0$ always hold, indicating that increases in both $S_{1}$ and $S_{2}$ help to reduce the loss. However, as shown intuitively in Fig. 1, there exists delicate balance between $S_{1}$ and $S_{2}$. When LR begins to anneal and $S_{2}$ starts to increase, the forward area $S_{1}$ of subsequent steps starts to diminish instead. Our equation aptly describes this delicate balance. In Sec. 4, we elaborate this topic in detail.
@@ 3.4 Experiments
Different warmup steps can result in different loss curves in training from scratch. During the warmup stage, neural networks are prone to random optimization, resulting in unpredictable outcomes (Hestness et al., 2017). Various studies, along with our own pilot experiments (Appendix A), show that LR warmup significantly accelerates model convergence. High gradient norms are usually observed during the LR warmup stage, especially in the initial steps of training (see Fig. 4). This indicates that model parameters undergo substantial updates during this stage. Therefore, in all our experiments, we linearly warmup LR to reach $\eta_{max}$ but compute $S_{1}$ and $S_{2}$ assuming a constant LR value $\eta_{max}$ in the warmup stage.
@@
We use standard experimental setups for LLM pre-training. In our main experiments, the training dataset is Fineweb (Penedo et al., 2024) and the validation dataset is RedPajama-CC (Computer, 2023). We train a $594$M non-embedding parameters LLAMA architecture-like model (Touvron et al., 2023) from scratch. We use AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The weight decay is set as $0.1$ and gradient clip is set as $1.0$. We set maximal learning rate as $2\times 10^{-4}$ and batch size as $4$M tokens. To verify the robustness of our formulation across different experimental settings, we have four other distinct experimental setups (see Appendix B). We adopt the decay factor or learning rate annealing $\lambda=0.999$ in our all experiments and discuss the impact of $\lambda$ in Sec. 6.1.
@@
Given a LRS of total step $s$, i.e., a learning rate value sequence $\{\eta_{1},\eta_{2}\,\cdots,\eta_{s}\}$, the value of $S_{1}$ and $S_{2}$ for each step can be computed using Eq. 1 in advance. To estimate $(L_{0},A,C,\alpha)$, we adopt a similar fitting method as used by Hoffmann et al. (2022). Specifically, we minimize the Huber loss (Huber, 1964) between the predicted and the observed log loss values using the L-BFGS algorithm (Nocedal, 1980):
$\min_{L_{0},A,C,\alpha}\quad\sum_{\mathrm{Step~{}}i}\mathrm{Huber}_{\delta}\left(\log\hat{L}(i)-\log L(i)\right).$ (6)
We use the implementation of the minimize method provided by the scipy library. Huber loss is to enhance the robustness of the fitting process and we set the $\delta$ value of Huber loss to $1.0\times 10^{-3}$. To address the potential issue of local minima during fitting, we select the optimal fit by testing across a range of initial conditions. Note that in practice, we can also fit the loss curves generated by multiple LRS using the same set of parameters $(L_{0},A,C,\alpha)$. In this situation, we sum the Huber losses from Eq. 6 of all fitted LRS.
@@
We fit Eq. 1 on the loss curves under constant and cosine LRS with 20K total steps (see Fig. 2), and then predict the full loss curves under several unseen LRS with 60K total steps (see Fig. 3). The results show an almost perfect fit, achieving a coefficient of determination ($R^{2}$) greater than $0.999$. This underscores the robust capability of our equation to accurately fit loss curves across diverse LRS using a single parameter tuple.
The prediction results in Fig. 3 indicate that our formulation is broadly applicable and generalizes robustly across four unseen LRS, with a mean prediction error as low as 0.2%. Moreover, our equation can accurately predict losses even for complex LRS that include multiple LR re-warmup stages (Fig. 3(e)), despite that the loss curves used for fitting do not contain any LR re-warmup stages.
Extensive Experiments on Different Setups.
To demonstrate the broad applicability of our proposed equation, we conduct additional fitting and prediction experiments using various setups. (1) We use an alternative set of training hyper-parameters (Appendix C.1); (2) We test our equation on the Mixture of Experts (MoE) architecture (Appendix C.2); (3) We apply our equation to predict loss curves for a much longer (10×10$\times$10 ×) training run involving a 1.7B parameter model trained on 1.4T tokens (Appendix C.3). (4) We fit the loss curves of open-sourced models, including BLOOM-176B trained on 300B tokens (BigScience, 2022) and OLMo-1B trained on 2T tokens  (Groeneveld et al., 2024) (Appendix C.4). All experiments produce excellent results, indicating that our equation is effective across diverse experimental setups, including different training hyper-parameters, architecture, model sizes, and dataset scales.
(a) The loss drop brought by LR annealing for different model size $N$. Dashed lines represent trends over steps. The loss drop brought by LR annealing scales with data size and model size.
(b) Curve fitting and prediction on cosine LRS (60K steps to $\eta_{min}=0.1\cdot\eta_{max}$) of different model sizes using our $N$-extended scaling law. Results for N=1214.4M are predicted.
Figure 7: The loss drop brought by LR annealing (left) and the $N$-extended full loss curve fitting and prediction (right). It suggests that $S_{2}\propto N^{\gamma}$ is reasonable from both experimental phenomena and accurate curve prediction.
@@ 3.5 Extension to Model Size Scaling
Loss Drop During Annealing Scales with Model Size $N$.
We explore the effect of model size $N$ on the loss drop during the annealing stage. Specifically, we compare the final losses obtained with a constant LRS and a WSD LRS (10% cosine annealing to $\eta_{min}=0$) to estimate the loss drop due to LR annealing. We conduct this experiment on different total steps and different model sizes. The experimental results are shown in Fig. 7(a). It suggests that the loss drop from LR annealing scales with both annealing steps and model sizes. This implies that the annealing area $S_{2}$ in our equation should also increase as the model size $N$ increases. We suppose there is a simple relationship of $S_{2}\propto N^{\gamma}$ where $\gamma$ is a positive constant.
Model Size Scaling.
Building on the experiments and analysis above, we extend our proposed Eq. 1 to incorporate model size scaling, based on traditional scaling laws (Eq. 2):
$L(s,N)=L_{0}+A\cdot S_{1}^{-\alpha}+B\cdot N^{-\beta}-C\cdot S_{2}\cdot N^{\gamma},$ (7)
where $N$ is the number of non-embedding model parameters, and $B$, $\beta$, $\gamma$ are positive constants related to $N$. We parameterize $S_{2}\propto N^{\gamma}$ via a multiplier $N^{\gamma}$ to the original annealing term $-C\cdot S_{2}$.
Fitting and Prediction with Model Size.
We validate Eq. 7 by fitting the full loss curves of models with varying sizes. We then apply the obtained equation to predict full loss curve on the unseen largest model size. Results in Fig. 7(b) show an almost perfect fit ($R^{2}>0.998$) and prediction for entire training dynamics of larger-scale models. This indicates the effectiveness and robustness of our proposed $N$-extended equation. Additional $N$-extended experiments with other setups further confirm the robustness of our formulation (see detail in Appendix C.5).
@@ 4 Takeaways: Experimental Findings Verification and Explanation
We apply our proposed formulation to validate and provide a theoretical explanation for numerous existing experimental findings regarding the training dynamics of language models. These key insights also guide researchers in selecting critical LRS before initiating model training. An interesting summary is that
@@
Our equation helps researchers understand why loss values drop more sharply when LR anneals. This phenomenon has been widely observed in many previous studies. Without loss of generality, consider the fitted equation from Fig. 2. We analyze how the $S_{1}$-item ($A\cdot S_{1}^{-\alpha}$) and the negative $S_{2}$-item ($-C\cdot S_{2}$) impact the loss values (notated as “$\Delta$Loss”) obtained across a WSD scheduler (see Fig. 8). It can be seen that in the annealing stage, the negative $S_{2}$-item has a much more significant impact on the overall loss than the $S_{1}$-item. Therefore, the loss drops more sharply compared to the stage with constant LR. In conclusion, LR annealing increases the annealing area, which further leads to a drastic decrease for the validation loss.
Figure 8: How S1-item and negative $S_{2}$-item changes in a WSD LRS. The upper and lower shaded area indicate the loss drop brought by $S_{1}$ and $S_{2}$, respectively, in the annealing stage.
4.2 Determining cosine cycle length and minimum LR in cosine LRS.
Many papers have found that in LLM pre-training using cosine LRS, setting the cosine cycle length $T$ as the total steps $S$, and setting min LR as nearly 0 (rather than 10% max LR) can lead to the optimal loss (Hoffmann et al., 2022; Hu et al., 2024; Hägele et al., 2024; Parmar et al., 2024). We theoretically validate this observation using our equation in Fig. 9. The predicted loss curve with $T=S$ and a minimum LR of 0 indeed achieves the optimal loss in the final step. Moreover, our equation gives a quite intuitive explanation: setting $T>S$ leads to incomplete annealing, while $T<S$ leads to a small forward area $S_{1}$ due to early annealing. Thus, the optimal configuration is to set $T$ equal to $S$. Also, setting the minimum LR to 00 maximizes the annealing amount, thereby increasing the annealing area $S_{2}$, which facilitates lower final loss.
Figure 9: Predicted loss curves of different cycle length $T$ and min LR in cosine LRS. The results well align with previous studies.
@@
(a) The predicted loss curve of 10K total steps.
(b) The predicted loss curve of 100K total steps.
@@
Figure 10: Predicted loss curves and final losses for the constant and cosine LRS.
In our experiments shown in Fig. 2, we observe that the constant LRS can sometimes yield lower final loss values than the cosine LRS. To validate this phenomenon, we use our equation to predict the loss curve of a constant and cosine LRS for 10K and 100K total steps in Fig. 10. It can be seen that with more training steps, the cosine LRS outperform the constant LRS, while with less training steps, the constant LRS is better. Moreover, Fig. 10(c) shows the predicted final loss of different total steps using constant and cosine LRS. It further convincingly suggests that constant LRS indeed gets a lower loss if setting small total steps, but the scaling slope is smaller than cosine LRS’s, resulting in higher loss in more steps.
We can further analyze this phenomenon by taking the derivative of $S_{1}$ and $S_{2}$ of Eq. 1. Specifically, $|\frac{\partial L}{\partial S_{1}}|=\alpha A\cdot S_{1}^{-\alpha-1}$ is a power-law decreasing function while $|\frac{\partial L}{\partial S_{2}}|=C$ is a constant. When the training step $s$ is small, $|\frac{\partial L}{\partial S_{1}}|$ is larger than $|\frac{\partial L}{\partial S_{2}}|$ since we have a small $S_{1}$. Therefore $S_{1}$ plays a dominant role over $S_{2}$. In this case, increasing $S_{1}$ by maintaining a large LR, i.e. using a constant LRS rather than a cosine LRS, leads to more sharply decrease of the loss value.
When the training step $s$ is large, $|\frac{\partial L}{\partial S_{1}}|$ becomes smaller than $|\frac{\partial L}{\partial S_{2}}|$. Therefore, $S_{2}$ plays a dominant role over $S_{1}$. In this case, increasing $S_{1}$ does not contribute much for decreasing the loss. It is time to start LR annealing to increase $S_{2}$. Interestingly, this formulation aligns with the idea of WSD LRS (Hu et al., 2024): In the early stages, the neural network is exploring globally and it is a suitable time to use a larger LR; In the later stages, the neural network is exploring locally and it is a suitable time to use a smaller LR. We will delve further into WSD LRS in the following subsections.